For numbers 8-15: An association of City Mayors conducted a study to determine the average number of times a family went to buy necessities in a week. They found that the mean is 4 times in a week. A random sample of 20 families were asked and found a mean of 5 times in a week and a standard deviation of 2. Use 5% significance level to test that the population mean is not equal to 5. Assume that the population is normally distributed.


1
Expert's answer
2021-06-28T03:32:25-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=5"

"H_1:\mu\not=5"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4-5}{2/\sqrt{20}}\approx-2.236"

Since it is observed that"|z|=2.236>1.96=z_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 5, at the "\alpha=0.05" significance level.

Using the P-value approach: The p-value is "p=2P(z<-2.236)=0.025347," and since "p=0.025347<0.05=\alpha," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 5, at the "\alpha=0.05" significance level.